# FOOL: Addressing the Downlink Bottleneck in Satellite Computing with Neural Feature Compression

### v2

@@ Abstract
This work presents FOOL, an OEC-native and task-agnostic feature compression method that preserves prediction performance. FOOL partitions high-resolution satellite imagery to maximize throughput. Further, it embeds context and leverages inter-tile dependencies to lower transfer costs with negligible overhead. While FOOL is a feature compressor, it can recover images with competitive scores on quality measures at lower bitrates. We extensively evaluate transfer cost reduction by including the peculiarity of intermittently available network connections in low earth orbit. Lastly, we test the feasibility of our system for standardized nanosatellite form factors. We demonstrate that FOOL permits downlinking over 100$\times$ the data volume without relying on prior information on the downstream tasks.
@@ 1 Introduction
As constellation sizes and sensor resolutions increase, downlink bandwidth cannot keep up with the accumulating data volume [7, 8]. Additional ground station equipment may prevent link saturation, building and maintaining them, including licensing the necessary frequencies, is a significant cost factor for satellite operation. As an alternative, Orbital Edge Computing (OEC) proposes processing data at the source [9, 10, 11, 12, 13]. Recent work on reducing bandwidth requirements in OEC is roughly categorizable in aggressive (task-oriented) filtering and compression [14]. The former relies on subjective value measures that restrict their practicability to coarse-grained tasks, such as de-duplication or cloud filtering. The latter constrains entire missions to particular tasks or prediction models. We argue that the limitations of existing compression or other data reduction approaches are particularly adverse to OEC.
@@
First, the CubeSat design is intended for short-duration missions [3] (typically up to 3-5 years), and despite waning prices, launching sensor networks in space is still associated with substantial logistical, administrative, and monetary costs. Therefore, it seems undesirable to designate entire constellations to a small subset of tasks and Deep Neural Network (DNN) architectures. More pressingly, irrespective of whether current codecs can prevent bottlenecks, they may undermine the effectiveness of entire missions. Precisely, the assumption of prediction models only requiring a subset of information for image reconstruction may lead to false confidence in a codec to reliably discern the salient signals. We argue the opposite holds, i.e., when the objective is to accommodate arbitrary downstream tasks with prediction models instead of human experts, there is less potential for rate reductions. Intuitively, two seemingly visually identical images may have subtle differences in pixel intensities, which a prediction model could leverage to overcome physiological restrictions.
@@
In summary, three conflicting objectives aggravate the challenges for OEC: (i) maximizing downlinking captures, (ii) ensuring the value of the captures by relying on as few assumptions on downstream tasks as possible, and (iii) minimizing the risk from unpredictable adverse effects on current and future prediction models. To this end, we propose drawing from recent work on neural feature compression with Shallow Variational Bottleneck Injection (SVBI) [15, 16, 17]. The idea of SVBI is to reduce discarding information necessary for arbitrary, practically relevant tasks by targeting the shallow representation of foundational models as a reconstruction target in the rate-distortion objective. In other words, rate reductions come from constraining the solution space with abstract high-level criteria rather than reifying target tasks with an explicit definition of value or expert-crafted labels. We investigate whether the SVBI framework is suitable for EO from a compression perspective and identify lower-level system considerations given the oppressive constraints of OEC. Then, we apply our insights to introduce a Tile Holistic Efficient Featured Oriented Orbital Learned (THE FOOL) compression method, which we will refer to as FOOL for short. FOOL alleviates the challenges of OEC by generalizing SVBI to improve compression performance while introducing more specific methods that aid in meeting the requirements of OEC and EO tasks. FOOL comprises a simple pipeline, a profiler, and a neural feature codec with a separate reconstruction model. The pipeline minimizes overhead by CPU-bound pre- and post-processing with concurrent task execution. The profiler identifies configurations that maximize data size reduction, factoring in intermittently available downlinks and the trade-off between processing throughput and lowering bitrate from more powerful but costlier transforms. The neural codec’s architecture includes task-agnostic context and synergizes with the profiler’s objective to maximize throughput with batch parallelization by exploiting inter-tile spatial dependencies.
@@ 2.1 Collaborative Inference and Data Compression
The Deep Learning aspect of our method draws from recent advancements in collaborative inference [19] and data compression [20]. The underlying compression algorithm and objective functions are derived and extended from previous publications such as Entropic Student (ES) [15, 16] and FrankenSplit (FS) [17]. Both works re-formulize the distortion term from lossy compression methods [21] and deploy lightweight models suitable for resource-constrained mobile devices. Besides introducing novel components to further lower transfer costs, FOOL considers the diverging requirements due to intrinsic differences between terrestrial and orbital remote sensing.
@@ 2.2 Preventing Link Saturation with Orbital Inference
The system aspect of our method aligns best with work focusing on getting the data to the ground for further processing instead of performing inference on board [22, 12, 9, 23]. We emphasize the high variability among fundamental design principles for OEC, as it is an emerging field, and a comprehensive literature review is not within the scope of this work. Instead, we discuss the approaches we found most promising and share some design elements with FOOL.
@@
Gadre et al. introduce Vista [24], a Joint Source Channel Coding (JSCC) system for LoRa-enabled CubeSats, designed to enhance low-latency downlink communication of satellite imagery and DNN inference. It shows significant improvements in image quality and classification performance through LoRa-channel-aware image encoding. Moreover, the evaluation assumes simple tasks not representative of practical EO. In contrast, FOOL decouples image recovery from the initial compression objective and ensures task-agnostic preservation of information.
@@
Lu et al. introduce STCOD [25], a JSCC system for efficient data transmission and object detection in optical remote sensing. STCOD integrates satellite computing to process images in space, distinguishing between regions of interest (ROIs) and backgrounds. It shows promising results with a block-based adaptive sampling method, prioritizing transmitting valuable image blocks using fountain code [26]. The caveat is that ROI detectors that can reliably prevent predictive loss for downstream tasks require strong biases regarding sensor properties and task granularity. FOOL includes task-agnostic context, with significantly less overhead and more robustness towards varying conditions than an ROI detector. Furthermore, it is end-to-end optimized with the other compression model components without relying on the same biases or expert-crafted labels.
@@
Thematically, our work resembles Kodan by Denby et al. [27] the closest. Like FOOL, Kodan treats channel conditions as an orthogonal problem and primarily focuses on source coding to address the downlink and computational bottlenecks. Kodan uses a reference application for satellite data analysis and a representative dataset to create specialized small models. Once in orbit, it dynamically selects the best models for each data sample to maximize the value of data transmitted within computational limitations. Kodan’s excellent system design is promising but relies on assumptions that hinder practicability and the potential for meaningful rate reductions. Conversely, FOOL follows a different design philosophy by treating the downlink bottleneck primarily as a compression problem. We do not treat the computational deadline as a hard temporal constraint to decouple the method to a particular system design. FOOL’s profiler measures key performance indicators on the pixel level. Given hardware limitations, the aim is to reduce transfer cost by balancing the lower bitrate of more powerful encoders and the gain in processing throughput of more lightweight encoders.
@@ 3.1 The Downlink Bottleneck
There are two overarching objectives for a codec and the system in which we deploy its encoder. The system’s objective is to process and encode large volumes of high-dimensional data, given the physical limitations of LEO (nano-) satellites. A 3U nanosatellite following the CubeSat standard is limited to 10cm×$\times$10cm×$\times$30cm and 4kg [30] with restricted power supply by using solar harvesting [31]. The compression objective is to achieve a sufficiently low bitrate while maintaining the data’s integrity. The following elaborates on the challenges of conceiving a method that fulfills our criteria and the limitations of applying existing codecs.
@@ 3.2 Limitations of Codecs
Given remote image captures and a set of unknown associated object detection tasks, we seek a transformation of the captures into representations that minimize transfer costs and loss of information that may impact any detection tasks. We refer to generalizability as a measure of how well a method can minimize the predictive loss on unknown detection tasks. For example, a purely task-oriented encoding (e.g., [32]) can retain information for a set of explicitly defined tasks. Still, it does not generalize as the transformed data is unusable for non-overlapping tasks Besides bent pipes, textitlossless codecs are the only approach with easily understood guarantees on generalization. Nevertheless, lossless compression cannot adequately address the downlink bottleneck due to theoretical lower bounds. Promising alternatives are lossy methods that relax the requirement of relying on identical reconstruction for generalization. More formally, given a distortion measure $\mathcal{D}$, a constraint $D_{c}$ bounds the minimal bitrate to [33]:
@@ 3.3 Shallow Variational Bottleneck Injection
SVBI trains neural codecs by replacing the distortion term of the rate-distortion objective of variational image compression models [34] with head distillation (HD) [43, 44, 15, 16, 17]. Figure 2 illustrates an example, where the HD distortion measure penalizes a compression model for not sufficiently approximating the shallow representation of a pre-trained foundational model. We define a foundational model as a pre-trained DNN that can accommodate multiple tasks by attaching predictors or fine-tuning the deeper layers. In Knowledge Distillation (KD) terminology, the codec is referred to as the student and the shallow layers of a foundational model as the teacher. Note that KD is not this work’s focus, as HD diverges from the typical KD objective. The intuition behind SVBI is that if a codec can reconstruct the representation of a foundational model, then the representation is sufficient for at least all tasks associated with that model.
@@ 3.3.1 The Effectiveness of Shallow Features
First, the features are not human-readable, i.e., we cannot overlay the bounding boxes on the images. We could infer the global coordinates to present boxed overlaid on previously captured satellite imagery. This may suffice for observing (semi-) permanent objects (e.g., landmarks) but certainly not for ephemeral or moving objects (e.g., tracking the movement of vessels). Second, even if the trend toward transfer learning with foundational models [45, 46] can accommodate various predictors, client preferences for architectures may vary.
@@ 3.3.2 Rate Reductions by Task Specificity
The idea of task-oriented communication is that messages for model prediction may require less information than human domain experts, i.e., that it should be possible to reduce bitrate by not (exclusively) using input reconstruction as the distortion measure. We argue that this assumption contradicts empirical evidence demonstrating models outperforming human experts in various image-related tasks, i.e., machines can detect signals and patterns that humans physiologically or intellectually cannot. Rather, the opposite should hold, i.e., when compressing for quality using domain experts as judges, we should see more potential for rate savings, not less. The claim is consistent with the results in Figure 1 where codecs with high reconstruction quality result in images that are deceptively similar to the input (details in Section 6.3). Conclusively, rate reductions are from task specificity of the distortion measure, irrespective of the input interface, whether it is a particular layer of a DNN architecture, human receptors, or textual encoding. Note that this hols, even if when limiting measures to discriminative task objectives without any image reconstruction. Besides visualizing Equation 8, the input image illustrates a practical example. The frog subset of ImageNet distinguishes between Tree Frogs, Bullfrogs, and Tailed Frogs. Since these frog species have distinct figures and dominant colors, the more delicate characteristics of a tree frog are redundant for ImageNet classification. The network gradually discards information regarding the fine-grained blue-yellow colored patterns, permitting only the recovery of general shape and environment from the deep features. The deeper the features, the less structure and detail are present, which may be redundant for the task. Now, suppose training a codec where the encoder retains the minimal information necessary to reconstruct the output of the deepest layers for a classification task (e.g., similar to Vista [24]). Then, we can reduce the transfer cost to as low as $\log_{2}{({\#\text{labels}})}$ without predictive loss. However, we may lack the information for other tasks, i.e., there is a trade-off between generalization and the lower bound on the bitrate. In contrast, targeting shallow features for compression may strike a balance between aiming to retain information for all possible downstream tasks and only emphasizing the salient regions for the tasks associated with a foundational model. Arguably, the limitation is negligible, as maintainers will train foundational models with useful tasks in mind.
@@ 4.1 Compression and Prediction Request Flow
For samples processed by FOOL, there is a single encoder. The output $\boldsymbol{\hat{{\mathrm{y}}}}$ is forwarded to the detection pipeline, skipping the shallow layers. A decoder transforms $\boldsymbol{\hat{{\mathrm{y}}}}$ into a target representation. A predictor outputs bounding boxes for a specified task. An image reconstruction model optionally restores the latent to a human-interpretable image to overlay the bounding boxes. Samples downlinked with bent pipe or some image codec are forwarded to the shallow layers, skipping the FOOL decoder and reconstruction model. This section focuses on the pipeline, before Section 5 introduces the compression method.
@@ 5.1 Model Building Blocks
Figure 6 illustrates the architecture’s building blocks and how it organizes the primary networks for transform coding (Section 3.2). The primary networks have four stages that control the depth and width. Each stage has at most one lightweight attention block and at least one residual block. A residual block optionally up or downsamples the spatial dimensions. A stage’s width and depth parameters configure the number of channels and residual blocks within a stage.

### v3

@@ Abstract
This work presents an OEC-native and task-agnostic feature compression method that preserves prediction performance and partitions high-resolution satellite imagery to maximize throughput. Further, it embeds context and leverages inter-tile dependencies to lower transfer costs with negligible overhead. While the encoding prioritizes features for downstream tasks, we can reliably recover images with competitive scores on quality measures at lower bitrates. We extensively evaluate transfer cost reduction by including the peculiarity of intermittently available network connections in low earth orbit. Lastly, we test the feasibility of our system for standardized nanosatellite form factors. We demonstrate that the proposed approach permits downlinking over 100$\times$ the data volume without relying on prior information on the downstream tasks.
@@ 1 Introduction
As constellation sizes and sensor resolutions increase, downlink bandwidth cannot keep up with the accumulating data volume [7, 8]. Additional ground station equipment may prevent link saturation. However, building and maintaining them, including licensing the necessary frequencies, is a significant cost factor for satellite operation. As an alternative, Orbital Edge Computing (OEC) proposes processing data at the source [9, 10, 11, 12, 13]. Recent work on reducing bandwidth requirements in OEC is roughly categorizable in aggressive (task-oriented) filtering and compression [14]. The former relies on subjective value measures that restrict their practicability to coarse-grained tasks, such as de-duplication or cloud filtering. The latter constrains entire missions to particular tasks or prediction models. We argue that the limitations of existing compression or other data reduction approaches are particularly adverse to OEC.
@@
First, the CubeSat design is intended for short-duration missions [3] (typically up to 3-5 years), and despite waning prices, launching sensor networks in space is still associated with substantial logistical, administrative, and monetary costs. Therefore, it seems undesirable to designate entire constellations to a small subset of tasks and Deep Neural Network (DNN) architectures. More pressingly, irrespective of whether current codecs can prevent bottlenecks, they may undermine the effectiveness of entire missions. Precisely, the assumption that prediction models only require a subset of information for image reconstruction may lead to false confidence in a codec to reliably discern the salient signals. We argue the opposite holds, i.e., when the objective is to accommodate arbitrary downstream tasks with prediction models instead of human experts, there is less potential for rate reductions. Intuitively, two seemingly visually identical images may have subtle differences in pixel intensities, which a prediction model could leverage to overcome physiological restrictions.
@@
In summary, three conflicting objectives aggravate the challenges for OEC: (i) maximizing downlinking captures, (ii) ensuring the value of the captures by relying on as few assumptions on downstream tasks as possible, and (iii) minimizing the risk from unpredictable adverse effects on current and future prediction models. To this end, we propose drawing from recent work on neural feature compression with Shallow Variational Bottleneck Injection (SVBI) [15, 16, 17]. The idea of SVBI is to reduce discarding information necessary for arbitrary, practically relevant tasks by targeting the shallow representation of foundational models as a reconstruction target in the rate-distortion objective. In other words, rate reductions come from constraining the solution space with abstract high-level criteria rather than reifying target tasks with an explicit definition of value or expert-crafted labels. We investigate whether the SVBI framework is suitable for EO from a compression perspective and identify lower-level system considerations given the oppressive constraints of OEC. Then, we apply our insights to introduce a Tile Holistic Efficient Featured Oriented Orbital Learned (THE FOOL) compression method, which we will refer to as FOOL for short. FOOL alleviates the challenges of OEC by generalizing SVBI to improve compression performance while introducing more specific methods that aid in meeting the requirements of OEC and EO tasks. FOOL comprises a profiler, a neural feature codec with a separate reconstruction model, and a simple pipeline. The profiler identifies configurations that maximize data size reduction, factoring in intermittently available downlinks and the trade-off between processing throughput and lowering bitrate from more powerful but costlier transforms. The neural codec’s architecture includes task-agnostic context and synergizes with the profiler’s objective to maximize throughput with batch parallelization by exploiting inter-tile spatial dependencies. The pipeline minimizes overhead by CPU-bound pre- and post-processing with concurrent task execution.
@@ 2.1 Collaborative Inference and Data Compression
The Deep Learning aspect of our method draws from recent advancements in collaborative inference [19] and data compression [20]. The underlying compression algorithm and objective functions are derived and extended from our previous work [17], which re-formulizes the distortion term from lossy compression methods [21] and deploys lightweight models suitable for resource-constrained mobile devices. Besides introducing novel components to further lower transfer costs, FOOL considers the diverging requirements due to intrinsic differences between terrestrial and orbital remote sensing.
@@ 2.2 Preventing Link Saturation with Orbital Inference
The system aspect of our method aligns best with work focusing on getting the data to the ground for further processing instead of performing inference on board [22, 12, 9, 23]. We emphasize the high variability among fundamental design principles for OEC, as it is an emerging field, and a comprehensive literature review is not within the scope of this work. In summary, we found that current approaches focus on designing complex systems tailored to specific conditions and rely on strong assumptions limiting their applicability. Moreover, they may adequately model the system conditions but run experiments on toy tasks or on low-resolution images. Contrastingly, FOOL is a holistic approach to the downlink problem that considers satellite systems and imagery properties. The following discusses the approaches we find most promising as representatives in their general direction.
@@
Gadre et al. introduce Vista [24], a Joint Source Channel Coding (JSCC) system for LoRa-enabled CubeSats designed to enhance low-latency downlink communication of satellite imagery and DNN inference. It shows significant improvements in image quality and classification performance through LoRa-channel-aware image encoding. Moreover, the evaluation assumes simple tasks that are not representative of practical EO. In contrast, FOOL decouples image recovery from the initial compression objective and ensures task-agnostic preservation of information.
@@
Lu et al. introduce STCOD [25], a JSCC system for efficient data transmission and object detection in optical remote sensing. STCOD integrates satellite computing to process images in space, distinguishing between regions of interest (ROIs) and backgrounds. It shows promising results with a block-based adaptive sampling method, prioritizing transmitting valuable image blocks using fountain code [26]. The caveat is that ROI detectors that can reliably prevent predictive loss for downstream tasks require strong biases regarding sensor and task properties. FOOL includes task-agnostic context, with significantly less overhead and more robustness towards varying conditions than an ROI detector. Furthermore, it is end-to-end optimized with the other compression model components without relying on the same biases or expert-crafted labels.
@@
Thematically, our work resembles Kodan by Denby et al. [27] the closest. Like FOOL, Kodan treats channel conditions as an orthogonal problem and primarily focuses on source coding to address the downlink and computational bottlenecks. Kodan uses a reference application for satellite data analysis and a representative dataset to create specialized small models. Once in orbit, it dynamically selects the best models for each data sample to maximize the value of data transmitted within computational limitations. Kodan’s excellent system design is promising but relies on assumptions that hinder practicability and the potential for meaningful rate reductions. Unlike Kodan, we follow a different design philosophy by treating the downlink bottleneck primarily as a compression problem. Further, we do not treat the computational deadline as a hard temporal constraint to decouple the method to a particular system design, as reflected by FOOL’s profiler measuring key performance indicators on the pixel level. Given hardware limitations, the aim is to reduce transfer costs by balancing the lower bitrate of more powerful encoders and the gain in processing throughput of more lightweight encoders.
@@ 3.1 The Downlink Bottleneck
There are two overarching objectives for a codec and the system we deploy its encoder. The system’s objective is to process and encode large volumes of high-dimensional data, given the physical limitations of LEO (nano-) satellites. A 3U nanosatellite following the CubeSat standard is limited to 10cm×$\times$10cm×$\times$30cm and 4kg [30] with restricted power supply by using solar harvesting [31]. The compression objective is to achieve a sufficiently low bitrate while maintaining the data’s integrity. The following elaborates on the challenges of conceiving a method that fulfills our criteria and the limitations of applying existing codecs.
@@ 3.2 Limitations of Codecs
Given remote image captures and a set of unknown associated object detection tasks, we seek a transformation of the captures into representations that minimize transfer costs and loss of information that may impact any detection tasks. We refer to generalizability as a measure of how well a method can minimize the predictive loss on unknown detection tasks. For example, a purely task-oriented encoding (e.g., [32]) can retain information for a set of explicitly defined tasks. Still, it does not generalize as the transformed data is unusable for non-overlapping tasks Besides bent pipes, lossless codecs are the only approach with easily understood guarantees on generalization. Nevertheless, lossless compression cannot adequately address the downlink bottleneck due to theoretical lower bounds. Promising alternatives are lossy methods that relax the requirement of relying on identical reconstruction for generalization. More formally, given a distortion measure $\mathcal{D}$, a constraint $D_{c}$ bounds the minimal bitrate to [33]:
@@ 3.3 Shallow Variational Bottleneck Injection
SVBI trains neural codecs by replacing the distortion term of the rate-distortion objective of variational image compression models [34] with head distillation (HD) [43, 44, 15, 16, 17]. Figure 2 illustrates an example where the HD distortion measure penalizes a compression model for not sufficiently approximating the shallow representation of a pre-trained foundational model. We define a foundational model as a pre-trained DNN that can accommodate multiple tasks by attaching predictors or fine-tuning the deeper layers. In Knowledge Distillation (KD) terminology, the codec is referred to as the student and the shallow layers of a foundational model as the teacher. Note that KD is not this work’s focus, as HD diverges from the typical KD objective. The intuition behind SVBI is that if a codec can reconstruct the representation of a foundational model, then the representation is sufficient for at least all tasks associated with that model.
@@ 3.3.1 The Effectiveness of Shallow Features
First, the features are not human-readable, i.e., we cannot overlay the bounding boxes on the images. We could infer the global coordinates to present boxes overlaid on previously captured satellite imagery. This may suffice for observing (semi-) permanent objects (e.g., landmarks) but certainly not for ephemeral or moving objects (e.g., tracking the movement of vessels). Second, even if the trend toward transfer learning with foundational models [45, 46] can accommodate various predictors, client preferences for architectures may vary.
@@ 3.3.2 Rate Reductions by Task Specificity
The idea of task-oriented communication is that messages for model prediction may require less information than human domain experts, i.e., that it should be possible to reduce bitrate by not (exclusively) using input reconstruction as the distortion measure. We argue that this assumption contradicts empirical evidence demonstrating models outperforming human experts in various image-related tasks, i.e., machines can detect signals and patterns that humans physiologically or intellectually cannot. Rather, the opposite should hold, i.e., when compressing for quality using domain experts as judges, we should see more potential for rate savings, not less. The claim is consistent with the results in Figure 1 where codecs with high reconstruction quality result in images that are deceptively similar to the input (details in Section 6.3). Conclusively, rate reductions are from task specificity of the distortion measure, irrespective of the input interface, whether it is a particular layer of a DNN architecture, human receptors, or textual encoding. Note that this holds, even if when limiting measures to discriminative task objectives without any image reconstruction. Besides visualizing Equation 8, the input image illustrates a practical example. The frog subset of ImageNet distinguishes between Tree Frogs, Bullfrogs, and Tailed Frogs. Since these frog species have distinct figures and dominant colors, the more delicate characteristics of a tree frog are redundant for ImageNet classification. The network gradually discards information regarding the fine-grained blue-yellow colored patterns, permitting only the recovery of general shape and environment from the deep features. The deeper the features, the less structure and detail are present, which may be redundant for the task. Now, suppose training a codec where the encoder retains the minimal information necessary to reconstruct the output of the deepest layers for a classification task (e.g., similar to Vista [24]). Then, we can reduce the transfer cost to as low as $\log_{2}{({\#\text{labels}})}$ without predictive loss. However, we may lack the information for other tasks, i.e., there is a trade-off between generalization and the lower bound on the bitrate. In contrast, targeting shallow features for compression may strike a balance between aiming to retain information for all possible downstream tasks and only emphasizing the salient regions for the tasks associated with a foundational model. Arguably, the limitation is negligible, as maintainers will train foundational models with useful tasks in mind.
@@ 4.1 Compression and Prediction Request Flow
For samples processed by FOOL, there is a single encoder. The output $\boldsymbol{\hat{{\mathrm{y}}}}$ is forwarded to the detection pipeline, skipping the shallow layers. The detection pipeline for a single forward pass consists of a decoder, backbone and predictor. There may be multiple backbones the client can choose from, and each backbone may have multiple predictors. A decoder transforms $\boldsymbol{\hat{{\mathrm{y}}}}$ into an input representation for a particular backbone. A predictor outputs bounding boxes for a specified task. An image reconstruction model optionally restores the latent to a human-interpretable image to overlay the bounding boxes. Samples downlinked with bent pipe or some image codec are forwarded to the shallow layers, skipping the corresponding decoder and reconstruction model. This section focuses on the pipeline, before Section 5 introduces the compression method.
@@ 5.1 Model Building Blocks
Figure 6 illustrates the building blocks of the codec architecture we will introduce in Section 5.4 and how it organizes the primary networks for transform coding (Section 3.2). The primary networks have four stages that control the depth and width. Each stage has at most one lightweight attention block and at least one residual block. A residual block optionally up or downsamples the spatial dimensions. A stage’s width and depth parameters configure the number of channels and residual blocks within a stage.